OpenStudy (vera_ewing):

Solve for x, given the equation sqrt(x=9)-4=1

OpenStudy (vera_ewing):

@Directrix ?

OpenStudy (solomonzelman):

\(\large\color{black}{ \sqrt{x+9}-4=1 }\) add 4 to both sides.

OpenStudy (vera_ewing):

@SolomonZelman i'm pretty sure it's 16, solution is extraneous

OpenStudy (solomonzelman):

if you add 4 to both sides, raise both sides to the 2nd power, and then subtract 9 from both sides, you get the answer. it is not 16.

OpenStudy (solomonzelman):

it's not 16. can we do it together?

Directrix (directrix):

I'm thinking it is x = 34, not extraneous but the work will decide.

OpenStudy (solomonzelman):

\(\large\color{black}{ \sqrt{x+9}-4=1 }\) start from adding 4 to both sides, tell me, what do you get?

Directrix (directrix):